OpenStudy (anonymous):

among all pairs of numbers whos difference is 12, find a pair whose product is as small as possible. what is the minimum product?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

my bad 24

OpenStudy (kinggeorge):

Well, he have a number \(x\) and \(x+12\) so that \((x+12)-x=12\) Now we need to minimize \[x\cdot (x+12)=x^2+12x\]

OpenStudy (anonymous):

so its 24

OpenStudy (anonymous):

so the first part is 12 & the second part is 24?

OpenStudy (anonymous):

sure, I dunno. Im only in ALgebra 2

OpenStudy (kinggeorge):

There are two ways to do it. Option one: Do calculus. Option two: Since it's an upward opening parabola, we know it has a minimum at the vertex. We can find this vertex using the formula \[-{b \over 2a}\]Since \(a=1, b=12\), the vertex occurs at \[x=-{12 \over 2}=-6\]If we plug this value back into the equation, we find that \[y=(-6)^2+12(-6)=36-72=-24\]So the minimum product should be -36

OpenStudy (kinggeorge):

Namely, you get a product of \(-36\) and a difference of 12 when you use the numbers \(6\) and \(-6\).

OpenStudy (kinggeorge):

I made a typo in my explanation. When I said \[y=(−6)^2+12(−6)=36−72=−24\]It should have been\[y=(−6)^2+12(−6)=36−72=−36\]

OpenStudy (anonymous):

thank you soooooo much!!

OpenStudy (kinggeorge):

You're welcome.